If a hollow box of volume 40cm^3
is made to float on the water, then one-fourth of
its volume remains in the air. While floating a hole is made in the box so that
2.5 cm^3 water enters the box per second. What is the time required for the box to be
completely drowned.
1
Expert's answer
2018-10-15T15:50:09-0400

We have

V=V_0+dV/dt t

t=(V-V_0)/(dV/dt)=(1/4 V)/(dV/dt)

t=1/4 40/2.5=4 s.